OpenStudy (anonymous):

True or False? If x>0, the (ln x) ^6 = 6lnx. Counter example?

OpenStudy (anonymous):

False. It is only true if: \[\ln(x^6)=6\ln(x)\]

OpenStudy (jwt625):

agree

OpenStudy (anonymous):

so its only when the variable is raised to a higher power and not if the whole ln is raised to a power?

myininaya (myininaya):

counterexample: what if x=e

myininaya (myininaya):

\[(\ln(e))^6=1^6 \neq 6=6\ln(e)\]